OpenStudy (anonymous):

In radioactive decay series, what is the particle or element of each in of following? Describe types of radiation given off in this series. 1. 92238U  A + 24He 2. A  91234Pa + B 3. 91234Pa  C + –10e 4. C  90230Th + 24He 5. 90230Th  D + 24He 6. D  86222Rn + E 7. 86222Rn  84218Po + F 8. 84218Po  G + 24He 9. G  83214Bi + H 10. 83214Bi  I + –10e 11. I  J + 24He

thomaster (thomaster):

There's \(\alpha\)-decay. These are Helium particles \(\to \begin{matrix}4\\ 2\end{matrix}He\) Then you have \(\beta\)-decay. These are electrons \(\to \begin{matrix}~~~0\\ -1\end{matrix}e\) Don't know what particles in 2,6,7 and 9. You can write the nuclear equations easy in \(\LaTeX\): \(\begin{matrix}168\\ 77\end{matrix}Ir\to \begin{matrix}4\\ 2\end{matrix}He+\begin{matrix}164\\ 75\end{matrix}Re\) This is the code: `\(\begin{matrix}168\\ 77\end{matrix}Ir\to \begin{matrix}4\\ 2\end{matrix}He+\begin{matrix}164\\ 75\end{matrix}Re\)` \(\overline{\underline{\LARGE{\color{gold}{\star~}}\Large\tt\color{green}{I\;Hope\;this\;Helps!}\LARGE{\color{gold}{~\star}}}}\)